OpenStudy (anonymous):

What is the integral of e^-4x?

OpenStudy (anonymous):

did u mean \[\int e^{-4x}dx\]

OpenStudy (anonymous):

in this case, you can set \(u=-4x\). \(du=-4dx\) \[dx=-\frac{du}{4}\] hence, your question become: \[-\frac{1}{4}\int e^u du = -\frac{1}{4}e^u+C = -\frac{1}{4}e^{-4x}+C\]